The area of a rectangle of length x is given by 3x² +5x. Find the width of the rectangle
Answer by ptaylor(1894)


Let W=width of rectangle
Now we know that the area of a rectangle =Length times Width. So:
3x^2+5x=x*W we can rewrite this as follows:
x(3x+5)=x*W divide both sides by x (x cannot be equal to zero)
W=3x+5-----------------------ans
CK
x(3x+5)=3x^2+5x
3x^2+5x=3x^2+5x

Hope this helps---ptaylor